# Cooperative and Asynchronous Transformer-based Mission Planning for Heterogeneous Teams of Mobile Robots

### v1

@@ VII Conclusions & Future Work
This paper introduced CATMiP, a novel framework for coordinating heterogeneous multi-robot teams in environments with communication constraints. The proposed CMD-POMDP model provides the mathematical foundations of asynchronous and decentralized decision-making of heterogeneous agents by incorporating class-based distinctions across its components. Leveraging the transformer’s ability to handle variable input sequence lengths, the Multi-Agent Transformer architecture was extended to develop scalable coordination strategies that can be executed in a distributed manner by any number of agents.
@@
CATMiP demonstrated robustness and scalability across different team sizes, compositions, and environmental complexities in simulations of search and target acquisition tasks. The results highlighted the framework’s adaptability to sporadic communication, asynchronous operations, and varied map conditions, achieving high mission success rates even under strict communication constraints. By addressing key challenges such as communication dropout, asynchronous operations, and agent heterogeneity, CATMiP shows significant potential for real-world applications where mobile robots must operate under resource limitations and unpredictable conditions.
@@
Future research will investigate integrating temporal memory into the network, expanding the framework to scenarios involving dynamic targets. The system’s performance will also be evaluated in more realistic 3D simulation environments as well as real-world experiments.

### v2

@@ 8 Conclusions & Future Work
This paper introduced CATMiP, a novel framework for coordinating heterogeneous multi-robot teams in environments with communication constraints. The proposed CMacDec-POMDP model provides the mathematical foundations of asynchronous and decentralized decision-making of heterogeneous agents by incorporating class-based distinctions across its components. Leveraging the transformer’s ability to handle variable input sequence lengths, the Multi-Agent Transformer architecture was extended to develop scalable coordination strategies that can be executed in a distributed manner by any number of agents.
@@
CATMiP demonstrated robustness and scalability across different team sizes, compositions, and environmental complexities in simulations of search and target acquisition tasks. The results highlighted the framework’s adaptability to sporadic communication, asynchronous operations, and varied map conditions, achieving high mission success rates and competitive performance even under strict communication constraints. By addressing key challenges such as communication dropout, asynchronous operations, and agent heterogeneity, CATMiP shows significant potential for real-world applications where different types of mobile robots must cooperate under resource limitations and unpredictable conditions.
@@
Future research will investigate integrating temporal memory into the network and expanding the framework to scenarios involving dynamic targets. The system’s performance will also be evaluated in more realistic 3D simulation environments as well as real-world experiments.